OpenStudy (anonymous):

if i have x-2 over 3x+9 times 2x+6 over 2x-4 , how do i solve that ?

OpenStudy (anonymous):

\[\frac{x-2}{3x+9}*\frac{2x+6}{2x-4}\]

OpenStudy (anonymous):

yes how do i solve that ?

OpenStudy (phi):

do you know that you can do this: (3 x + 3*3) = 3(x+3) See if you make progress using this idea.